Logiweb dictionary of clueless

Up Help

0 0 clueless

1 0 Weakening 2 0 ConditionedMP 3 0 PermuteAntecedents 4 0 ImplyTransitivity 5 0 EqualReflexivity 6 0 EqualSymmetry 7 0 EqualTransitivity 8 0 Mendelson32d 9 0 substantiation 10 0 Mendelson32fBase 11 0 Mendelson32fInduction 12 0 Mendelson32f 13 0 Mendelson32gBase 14 0 Mendelson32gInduction 15 0 Mendelson32g 16 0 Switch 17 0 AddOne 18 0 Mendelson32dRule 19 0 Mendelson32dConditionedRule 20 0 EqualTransitivityRule 21 0 EqualTransitivityConditionedRule 22 0 PlusCommutativityBase 23 0 PlusCommutativityInduction 24 0 PlusCommutativity 25 0 DoubleMP 26 0 InductionMacro 27 0 EqualSymmetryRule 28 0 DoubleConditionedMP The pyk compiler, version 0.grue.20050603 by Klaus Grue GRD-2005-06-21.UTC:18:57:23.412938 = MJD-53542.TAI:18:57:55.412938 = LGT-4626097075412938e-6